Bwindi Impenetrable Forest (10.01.2012 to 12.31.2012)

You are here

PDF versionPDF version
October - December
Communications activities: 

No communication activities at the site

Unusual Events at the TEAM Site: 

No unusual events at the TEAM site

New Species: 

No new species identified at the site

Protocol Activities: 

Vegetation plots

Measurement error estimation for the diameter takers was conducted. A refresher training for all the crew was also conducted before plot re-measurement. Three of the six vegetation plots were re-measured. New recruits were identified, tagged and mapped.

Climate station:

Climate data for this quarter (with exception of December) has been uploaded onto the TEAM portal. Bi-weekly maintenance of the climate station has been carried out in accordance with the protocol. The maintenance log has also been uploaded onto the portal. Issues with temperature/humidity sensor 2 were reported to the technical director.

Camera traps:

Images from the camera traps were identified using the DeskTEAM software. Data DVDs were sent to the TEAM Network secretariat.

Protocol Problems: 

No major problems to protocol implementation Issues with temperature/humidity sensor 2 were reported to the technical director.

Schedule Problems: 

No deviations from the proposed sampling schedule. all was on schedule.

Logistical Problems: 

Delay in funds to get to us makes planning a little tricky, otherwise, no major logistical problems encountered.


A technical workshop be included in the planning for the forthcoming TEAM Network meeting. The workshop will help to solve technical problems with site equipment e.g camera traps and weather station sensors.